40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Sep 19 01:49:52 2020, Koji, Update, General, M4.5 EQ in LA 
    Reply  Sat Sep 19 11:27:04 2020, rana, Update, General, M4.5 EQ in LA Sep18-EQ.pdf
       Reply  Sat Sep 19 18:08:34 2020, rana, Update, General, M4.5 EQ in LA 
    Reply  Sat Sep 19 19:14:59 2020, Koji, Update, General, ITMX released / ETMY UR magnet knocked off IMG_6218.jpeg
       Reply  Sun Sep 20 11:41:54 2020, gautam, Update, General, IMC re-locked 
       Reply  Mon Sep 21 12:40:58 2020, gautam, Update, General, Workable IFO recovery IFOrecovery.pngIMG_5345.JPG
Message ID: 15588     Entry time: Sun Sep 20 11:41:54 2020     In reply to: 15585
Author: gautam 
Type: Update 
Category: General 
Subject: IMC re-locked 

While I stopped by the lab this morning to pick up some things, I took the opportunity to continue the recovery.

  • IMC suspensions were sufficiently misaligned that the autolocker couldn't re-acqurie the lock. I manually recovered the alignment and now the IMC is locked again.
  • ETMY illuminator was left ON, I turned it off. In the process, I modified the illuminator ON/OFF script to be compatible with python3, but unfortunately, it was written in a way that doesn't permit backward compatibility, so now the illuminators can't be turned ON/OFF via the MEDM screen on pianosa (since the default python is 2.7 on that machine). But it does work on rossa, which I'm using as my primary workstation now (hence the change).
  • ITMX watchdog trip threshold was manually reset to the nominal value - the rampdown script was working, but the threshold was ~1400cts (normally ~200 cts) even at 1130am this morning (>12 hours after Koji's work yesterday evening), so I just accelerated the process.
  • Suspension realignment - using a mix of green and IR beams and the various cameras/photodiodes as diagnostics, I roughly restored the alignment of all the suspensions, except ETMY. I can see IR resonances in the X arm now.

At some point, we should run the suspension eigenmode routine (kick optics, let them ringdown, measure peak locations and Qs) to confirm that the remaining suspensions are okay, will also help in actuation re-allocation efforts on ETMY. But I didn't do this today.

Leaving the lab at 1150.

ELOG V3.1.3-